OpenStudy (anonymous):

Let ρAl represent the density of aluminum and ρFe that of iron. Find the radius of a solid aluminum sphere that balances a solid iron sphere of radius rFe on an equal-arm balance. (Use any variable or symbol stated above as necessary.)

OpenStudy (anonymous):

The volume of a sphere is 2/3 * PI * r^3 The mass of that sphere is denisty * volume = ρ * 2/3 * PI * r^3 If the masses of the two spheres are going to be equal... ρAl * 2/3 * PI * rAl^3 = ρFe * 2/3 * PI * rFe^3 And then cancel and re-arrange to find the answer.

OpenStudy (anonymous):

volume of sphere is 4/3pir^3

OpenStudy (anonymous):

You are quite right, vicky007. Don't know how that one slipped through the mental net. Well done. :-)